# PWDFT-SW: Extending the Limit of Plane-Wave DFT Calculations to 16K Atoms on the New Sunway Supercomputer

### v1

@@ I Introduction
Among the major applications in high-performance computing (HPC), the first-principles density functional theory (DFT) related applications are the most important ones in quantum mechanical material simulations [5]. The DFT calculation is essential in many fields because it can provide a direct link between atomic structure and corresponding physical properties [17]. Considering the gross importance of DFT calculations, there have been a number of works that aimed to develop accurate and efficient approaches for solving the Kohn-Sham DFT problem [31, 33, 25, 6, 15, 35, 29, 2]. There are several ways to numerically discretize the wavefunctions in DFT calculations to solve the Kohn-Sham DFT problem, such as numerical atomic orbitals (NAO) basis sets [16, 23, 20] and Gaussian-type orbital (GTO) basis sets [10]. However, a perceived drawback of these types of basis sets is that they are not systematically convergent for generic material systems. In contrast, the plane wave (PW) basis set [13, 12, 19] is preferred for its systematic convergence and high accuracy, even in strongly correlated physical systems. This preference stems from the PW basis set’s completeness, orthogonality, and nonlocality. These unique advantages have resulted in the arisen of PW-based DFT software in different fields, such as commercial software VASP [28, 14] and open source code Quantum Espresso [12, 11].
@@ II-A Background of PWDFT
Figure 1: Basic flowchart and 4 time-consuming steps with their corresponding numerical kernels (in the same background color) of PWDFT
Similar to many other DFT codes, PWDFT solves the eigenvalue problem in a self-consistent field (SCF) method. The SCF iterations in PWDFT contain two parts: (1) inner SCF for the eigenvalue problem of PW Hamiltonian and (2) outer SCF for checking the convergence of energy [32], as shown in Figure 1. At outer SCF, the Kohn-Sham equation $H\Psi_{i}=\varepsilon_{i}\Psi_{i}$ is solved to obtain the improvement of the wavefunction $\Psi_{i}$ for a given particle Hamiltonian $H_{i}$, and the wavefunction is expanded based on plane wave basis set. Outer SCF terminates when the output exchange energy $V_{xc}$ agrees to input wavefunctions. The inner SCF is mainly responsible for diagonalizing the PW Hamiltonian by an iterative conjugate gradient eigensolver PPCG or LOBPCG. The inner SCF begins when the Hamiltonian matrix is constructed by wavefunctions, which aim to solve eigenproblems and acquire eigenvalues and eigenfunctions. Then the charge density $\rho\left(r\right)$ is calculated by eigenfunctions according to $\rho\left(r\right)=\sum_{j=1}^{N_{e}}\left|\psi_{j}\right|^{2}$, where $\psi_{j}$ is the j-th eigenfunction and $N_{e}$ represents the total valence charge number. Then exchange-correlation potential $V_{xc}$ is calculated to be compared with the input for consistency. We present the major numerical kernels in each step on PWDFT on the right of Figure 1 to enhance the readability.
In theory, wave functions are constructed using an infinite plane wave basis. However, $E_{cut}$ is set to exclude waves with energy less than a certain input value. This cutoff energy determines the grid number in real space $N_{r}$ according to $\left(N_{r}\right)i=\sqrt{2E{\text{cut }}}L_{i}/\pi$, where $L_{i}$ is the length of the supercell in different dimensions (x, y, and z). In our experiments, the default value of $E_{cut}$ is set to 5 Hartree unless otherwise specified. Table I lists the grid number for certain systems in our experiments.
TABLE I: Grid Number of Typical Test Cases
@@ II-B Parallelization Scheme
There are three main parallelization schemes as shown in Figure 2 due to different intrinsic physical natures in PWDFT. The parallelization scheme is based on the distribution of the wavefunctions matrix. The first scheme is the row block partition. This scheme facilitates matrix-matrix multiplications such as $S=\Psi^{*}(H\Psi)$ when $H\Psi$ is obtained. The second scheme is the column block partition. In this scheme, each process holds some entire wavefunctions, allowing a single process to perform fast Fourier transformations (FFT) without any data transformation or communication. This data distribution pattern also simplifies the calculation of $H\Psi$, where $H$ is the Hamiltonian operator applied to wavefunctions. The third scheme is a 2-dimensional data partition used exclusively for Syevd, a ScaLAPACK function to acquire eigenvalues and eigenvectors. We perform MPI_Alltoallv and a ScaLAPACK routine Pdgemr2d to transform data patterns among the 3 schemes.
Figure 2: Parallel data and task distribution schemes
@@ II-C1 Limited Memory Size of Each Process
Figure 3 depicts the architecture of the SW26010Pro processor in the new Sunway supercomputer, comprising six core groups (CGs) each with one management processor element (MPE) and 64 computing processor elements (CPEs). Each processor is equipped with 96 GB of memory, evenly allocated at 16 GB per CG. The new Sunway supercomputer employs two memory partition schemes: (1) the common scheme, where each core group acts as a separate processor with 16 GB of memory, and (2) the large memory scheme, where six core groups in a process share 96 GB of memory and act as a single process. However, numerical libraries that are extensively used in PWDFT like FFTW [9], BLAS [8], and LAPACK [1] are tailed to the common scheme, implying that each process of PW-based DFT can only use the common memory partition scheme, which allocates a very limited 16 GB memory footprint per process.
Figure 3: Architecture of SW26010Pro Processor
The architecture of the many-core processor SW26010Pro on the new Sunway supercomputer is shown in Figure 3. It consists of 6 core groups (CGs), in which there is one management processor element (MPE) and $8×8$ computing processor elements (CPEs). The SW26010Pro processor has a memory capacity of 96 gigabytes (GB), which is divided equally among six core groups, with each group having 16 GB of memory. The new Sunway supercomputer employs two memory partition schemes for running applications: (1) the common scheme, where each core group acts as a separate processor with 16 GB of memory, and (2) the large memory scheme, where six core groups in a process share 96 GB of memory and act as a single process. Conventional DFT software relies heavily on numerical libraries such as FFTW [9], BLAS [8] and LAPACK [1] etc. However, these numerical libraries only support the common scheme, which implies that each process of PW-based DFT, distributed on each CG, can only allocate a very limited 16 GB memory footprint.
@@ III-A Refactor of the pseudopotential calculations
Figure 4: Refactor of pseudopotential process: (a) naive implementation where each process holds the entire pseudopotential, (b) the refactored approach where each process holds a portion of pseudopotential, in which the dotted, uncolored blocks indicate the absence of pseudopotential in the process.
The original procedure of applying pseudopotential to wave functions is illustrated in Figure 4 (a). These functions are distributed across various processes. For each wave function stored locally, the algorithm sequentially processes every atom, adjusting the wave function according to the corresponding pseudopotential. When the pseudopotential is distributed across processes, each process retains only a portion of the overall pseudopotential data. Therefore to apply the pseudopotential, it has to be retrieved from other processes through a network interface. In the case of a basic algorithm, this necessitates the redundant transmission of the same atomic pseudopotential across distinct wave functions.
Therefore, we apply the pseudopotential of an atom to all local wave functions once obtained, thereby eliminating redundant transmission of pseudopotential. Additionally, we employ a sliding window algorithm to address significant communication overhead, enabling overlapping computation and communication. This approach, illustrated in Figure 4 (b), allows a process to receive the upcoming pseudopotential within the sliding window while applying the current pseudopotential to the local wave functions. The communication overhead can be fully overlapped with the flexible window size.
@@ III-B2 Optimizations
We combine the data in the same column of each process into sub-blocks, as shown at the bottom of Figure 5. The number in each sub-block indicates the target position for the transformation. This sub-block method offers three benefits: (1) it increases the granularity of data transfer, (2) it significantly decreases the cost of the mapping matrix and (3) it enhances the temporal and spatial locality of the associated data transfer.
@@ III-C Communication Optimization
Figure 7: Performance of MPI_Allreduce regarding different communication volumes on the Sunway supercomputer
@@ III-C2 Algorithm and Implementation
Algorithm 1 Multistage Allreduce
@@ III-C3 Performance and Analysis
Figure 8: Allreduce performance speedup with the same communication volume and different processes
@@
Figure 9: Allreduce performance speedup with different communication volume and the same processes
@@ III-D Granularity-Aware Parallelization
TABLE II: Syevd time for Si8192 with different processes
@@
As discussed in Section III-C, MPI communication benefits from network topology and MPI implementation when the number of processes is an exponent of 2. This effect is even more pronounced in the Syevd routine, as we observe significant time-to-solution changes when the number of processes changes from 1000 to 1024, 2048 to 2056, and 4096 to 4112, as shown in Table II. Thus, there is a minimum threshold for the number of processes for a given atomic system, noted as $P_{min}$. We have developed a C++ library for memory analysis of PWDFT (discussed in Section 10), which allows us to analyze memory usage and determine the threshold for the number of processes. For a silicon atomic system, one atom corresponds to two wavefunctions, the minimum granularity for FFT and matrix multiplication. Therefore, the maximum number of processes is twice the number of atoms (noted as $A$).
@@ III-E SWUC-assisted Heterogeneous Multi-core Programming
Fortunately, a new programming framework called ShenWei Universal C/C++ (SWUC) [4] has been proposed to simplify programming with CPEs. Thanks to several new attributes and compiler directives, writing codes running on CPEs and MPE can be simply achieved by calling for lambda expressions, in which we divide data for each CPE and specify the codes to be executed. We adopt the framework in the calculation of pseudopotential, wavefunctions, ion density, and energy. We divide the data according to the layout of CPEs, batch data, and transfer it to the LDM of CPEs, and obtain the final results via lambda function.
However, for some more complex computing scenarios, SWUC is not applicable yet. In such cases, we have to use the Athread library, the most primitive interface to use CPEs, to accelerate PWDFT. The numerical result of the SWUPC-assisted CPE programming will be discussed in Section IV-D.
@@ III-F Other Optimization and Tool
In this subsection, we describe briefly our minor contributions to PWDFT on the new Sunway supercomputer.
Data Structure Optimization: When constructing the PW Hamiltonian, it’s necessary to look up the pseudopotential for a given atom. However, the original implementation of PWDFT, which targets general-purpose CPUs, used std::map from the C++ Standard Template Library to perform the look-up. This method suffers from poor performance due to its indirect access, especially when the system size increases. Additionally, the CPEs on SW26010Pro are unable to allocate from the main memory shared across cores, making std::map inaccessible on CPEs and hindering pseudopotential lookup acceleration. Considering that there are typically no more than 200 different types of atoms, to address these issues, we replace std::map with a small array. By indexing the array with the atom number, we are able to acquire the pseudopotential with one direct access, resulting in significant performance improvements, as shown in Figure 10. This new mechanism is also more CPE-friendly, as the LD Cache makes loading the pseudopotential efficient in case of direct indexing. Compared to the previous design, this method boosts the performance of the PW Hamiltonian step by a factor of 10.
Figure 10: Data Structure Optimization of Pseudopotential
@@ IV-A Setup of the Test Physical Systems and Testing Environment
Figure 11: Testing systems: (a) graphene system and (b) bulk silicon system
The testing systems include two parts: magic-angle twisted bilayer graphene (labeled as ”G”) and cubic silicon systems (labeled as ”Si”), as shown in Figure 11. These systems are denoted respectively as G11520, G16384, Si64, Si216, Si512, Si1000, Si4096, and Si8192. The corner marks indicate the number of atoms in each system.
In our experiment, the kinetic energy cutoff $E_{cut}$ is set to 5 Hatree unless otherwise specified. Our optimization approach has good effects for different values of $E_{cut}$. All numerical tests are performed on the new Sunway supercomputer, and each MPI process is applied to one CG (including 1 MPE and 64 CPEs) with separated 16 GB memory. We consider the time per outer SCF iteration as the time to solution.
@@ IV-B Strong Scalability
The speedup is compared with the minimum number of CGs (64 CGs for Si4096 and 512 CGs for Si8192) capable of performing the calculation. We use 64 to 2048 CGs when simulating Si4096 to test the strong scalability of PWDFT and 512 to 4096 CGs for Si8192. The results are shown in Figure 12. From the figure, we can see that, when scaling up to 2048 CGs (Si8192), PWDFT maintains a parallel efficiency of 48.5%, which is quite acceptable for PW-based DFT calculations. However, the parallel efficiency is not satisfactory when scaling up to half the number of atoms. Even so, we must note that such mediocre parallel efficiency is already much better than the original version and other PW-based DFT software. In Section III-D, due to the poor scalability of Syevd, the optimal process number is always selected in our granularity-aware parallelism for Syevd in the range of $P_{opt}\in[P_{min},2*A]\land P_{opt}=2^{k},k\in\mathbb{N}^{+}$. Hence, when the number of processes reaches the sufficient threshold to achieve the best performance for Syevd, it remains unchanged. Considering that Syevd often occupies a significant portion of the total computation time, it is reasonable to conclude that PWDFT does not exhibit excellent strong scalability and parallel efficiency. Additionally, massive global communication, as discussed in Section II-C2, also partially contributes to this inefficiency.
Figure 12: Strong scaling of PWDFT
@@
Figure 13: Strong scaling of PWDFT without syevd

### v2

@@ I Introduction
Among the major applications in high-performance computing (HPC), first-principles density functional theory (DFT) calculations hold significant importance for quantum mechanical simulations of materials [6]. DFT enables a direct connection between atomic structures and their associated physical properties, underpinning research in numerous scientific and engineering fields [22]. Given the critical role of DFT, extensive efforts have been made to develop efficient and accurate numerical approaches to solve the Kohn-Sham equations [40, 46, 30, 9, 20, 49, 35, 2]. Several numerical discretization methods exist for representing wavefunctions in DFT calculations, including numerical atomic orbital (NAO) basis sets [21, 28, 25] and Gaussian-type orbital (GTO) basis sets [13]. However, these localized orbital-based approaches are typically not systematically convergent for generic material systems. In contrast, the plane wave (PW) basis set [18, 16, 24] is widely favored due to its systematic convergence, accuracy, completeness, and orthogonality, making it particularly suitable even for strongly correlated systems. Consequently, PW-based DFT codes, such as the widely used commercial software VASP [34, 19] and the open-source package Quantum Espresso [16, 15], have gained extensive popularity in computational material science.
Recently, real-space discretization methods, such as SPARC [54] and DFT-FE [39], have demonstrated significant advantages in HPC environments due to their excellent parallel scalability and reduced communication overhead. These real-space methods avoid the global communication bottlenecks associated with the Fast Fourier Transform (FFT) operations required by PW methods. Nevertheless, the maturity, established software ecosystem, and robust convergence properties of PW-based methods continue to make them highly attractive and widely adopted in current research. In this work, we focus on the plane-wave-based PWDFT framework, aiming to exploit these advantages while enhancing computational efficiency and scalability.
@@ II-A Background of PWDFT
Figure 1: Basic flowchart and 4 time-consuming steps with their corresponding numerical kernels (in the same background color) of PWDFT.
Similar to many other DFT codes, PWDFT solves the eigenvalue problem in a self-consistent field (SCF) method. The whole SCF iterations in PWDFT contain two parts: (1) the inner eigenvalue problem of PW Hamiltonian and (2) SCF for checking the convergence of energy [43], as shown in Figure 1. SCF part, the Kohn-Sham equation $H\Psi_{i}=\varepsilon_{i}\Psi_{i}$ is solved to obtain the improvement of the wavefunction $\Psi_{i}$ for a given particle Hamiltonian $H_{i}$, and the wavefunction is expanded based on plane wave basis set. The SCF terminates when the local potential $V_{Hxc}$ agrees to the input wavefunctions. The inner eigenvalue problem is mainly responsible for diagonalizing the PW Hamiltonian by an iterative conjugate gradient eigensolver, PPCG or LOBPCG. The inner eigenvalue problem begins when the Hamiltonian matrix is constructed by wavefunctions, which aim to solve eigenproblems and acquire eigenvalues and eigenfunctions. Then the charge density $\rho\left(r\right)$ is calculated by eigenfunctions according to $\rho\left(r\right)=\sum_{j=1}^{N_{e}}\left|\psi_{j}\right|^{2}$, where $\psi_{j}$ is the j-th eigenfunction and $N_{e}$ represents the total valence charge number. Then exchange-correlation potential $V_{Hxc}$ is calculated to be compared with the input for consistency. We present the major numerical kernels in each step on PWDFT on the right of Figure 1 to enhance the readability.
In theory, wave functions are constructed using an infinite plane wave basis. However, $E_{cut}$ is set to include waves with energy less than a certain input value. This cutoff energy determines the grid number in real space $N_{r}$ according to $\left(N_{r}\right)_{i}=\sqrt{2E_{cut}}L_{i}/\pi$, where $L_{i}$ is the length of the supercell in different dimensions (x, y, and z). In our experiments, the default value of $E_{cut}$ is set to 5 Hartree unless otherwise specified. Table I lists the grid number for certain systems in our experiments.
TABLE I: Grid Number of Typical Test Cases.
@@ II-B Parallelization Scheme
There are three main parallelization schemes as shown in Figure 2 due to different intrinsic physical natures in PWDFT. The parallelization scheme is based on the distribution of the wavefunction matrix. The first scheme is the row block partition. This scheme facilitates matrix-matrix multiplications such as $S=\Psi^{*}(H\Psi)$ when $H\Psi$ is obtained. The second scheme is the column block partition. In this scheme, each process holds some entire wavefunctions, allowing a single process to perform a fast FFT without any data transformation or communication. This data distribution pattern also simplifies the calculation of $H\Psi$, where $H$ is the Hamiltonian operator applied to wavefunctions. The third scheme is a 2-dimensional data partition used exclusively for Syevd, a ScaLAPACK function to acquire eigenvalues and eigenvectors. We perform MPI_Alltoallv and a ScaLAPACK routine Pdgemr2d to transform data patterns among the 3 schemes.
Figure 2: Parallel data and task distribution schemes.
@@ II-C1 Limited Memory Size of Each Process
Figure 3 depicts the architecture of the SW26010Pro processor in the new Sunway supercomputer, comprising six core groups (CGs) each with one management processor element (MPE) and 64 computing processor elements (CPEs). Each processor is equipped with 96 GB of memory, evenly allocated at 16 GB per CG. The new Sunway supercomputer employs two memory partition schemes: (1) the common scheme, where each core group acts as a separate processor with 16 GB of memory, and (2) the large memory scheme, where six core groups in a process share 96 GB of memory and act as a single process. However, numerical libraries that are extensively used in PWDFT like FFTW [12], BLAS [11], and LAPACK [1] are tailored to the common scheme, implying that each process of PW-based DFT can only use the common memory partition scheme, which allocates a very limited 16 GB memory footprint per process.
Figure 3: Architecture of SW26010Pro Processor.
@@ III-A Refactor of the pseudopotential calculations
Figure 4: Refactor of pseudopotential process: (a) naive implementation where each process holds the entire pseudopotential, (b) the refactored approach where each process holds a portion of the pseudopotential, in which the dotted, uncolored blocks indicate the absence of pseudopotential in the process.
The original procedure of applying a pseudopotential to wave functions is illustrated in Figure 4 (a). These functions are distributed across various processes. For each wave function stored locally, the algorithm sequentially processes every atom, adjusting the wave function according to the corresponding pseudopotential. When the pseudopotential is distributed across processes, each process retains only a portion of the overall pseudopotential data. Therefore, to apply the pseudopotential, it has to be retrieved from other processes through a network interface. In the case of a basic algorithm, this necessitates the redundant transmission of the same atomic pseudopotential across distinct wave functions.
Therefore, we apply the pseudopotential of an atom to all local wave functions once obtained, thereby eliminating redundant transmission of the pseudopotential. Additionally, we employ a sliding window algorithm to address significant communication overhead, enabling overlapping computation and communication. This approach, illustrated in Figure 4 (b), allows a process to receive the upcoming pseudopotential within the sliding window while applying the current pseudopotential to the local wave functions. The communication overhead can be fully overlapped with the flexible window size.
@@ III-B2 Optimizations
We combine the data in the same column of each process into sub-blocks, as shown at the bottom of Figure 6. The number in each sub-block indicates the target position for the transformation. This sub-block method offers three benefits: (1) it increases the granularity of data transfer, (2) it significantly decreases the cost of the mapping matrix, and (3) it enhances the temporal and spatial locality of the associated data transfer.
@@
Figure 6: An example of the in-place data transformation. The circled numbers denote the procedures in the transformation.
@@ III-C Communication Optimization
Figure 7: Performance of MPI_Allreduce regarding different communication volumes on the Sunway supercomputer.
@@ III-C2 Algorithm and Implementation
Algorithm 1 Algorithm of Multistage Allreduce.
@@ III-C3 Performance and Analysis
Figure 8: Allreduce performance speedup with the same communication volume and different processes.
@@
Figure 9: Allreduce performance speedup with different communication volume and the same processes.
@@ III-D Granularity-Aware Parallelization
TABLE II: Syevd time for Si8192 with different processes.
@@
As discussed in Section III-C, MPI communication benefits from network topology and MPI implementation when the number of processes is a power of 2. This effect is even more pronounced in the Syevd routine, as we observe significant time-to-solution changes when the number of processes changes from 1000 to 1024, 2048 to 2056, and 4096 to 4112, as shown in Table II. Thus, there is a minimum threshold for the number of processes for a given atomic system, noted as $P_{min}$. We have developed a C++ library for memory analysis of PWDFT (discussed in Section 10), which allows us to analyze memory usage and determine the threshold for the number of processes. For a silicon atomic system, one atom corresponds to two wavefunctions, the minimum granularity for FFT and matrix multiplication. Therefore, the maximum number of processes is twice the number of atoms (noted as $A$).
@@ III-E SWUC-assisted Heterogeneous Multi-core Programming
Fortunately, a new programming framework called ShenWei Universal C/C++ (SWUC) [5] has been proposed to simplify programming with CPEs. Thanks to several new attributes and compiler directives, writing code running on CPEs and MPE can be simply achieved by calling lambda expressions, in which we divide data for each CPE and specify the code to be executed. We adopt the framework in the calculation of pseudopotential, wavefunctions, charge density, and energy. We divide the data according to the layout of CPEs, batch data, and transfer it to the LDM of CPEs, and obtain the final results via a lambda function.
However, for some more complex computing scenarios, SWUC is not yet applicable. In such cases, we have to use the Athread library, the most primitive interface to use CPEs, to accelerate PWDFT. The numerical result of the SWUPC-assisted CPE programming will be discussed in Section IV-D.
@@ III-F Other Optimization and Tool
In this subsection, we briefly describe our minor contributions to PWDFT on the new Sunway supercomputer.
Data Structure Optimization: When constructing the PW Hamiltonian, it’s necessary to look up the pseudopotential for a given atom. However, the original implementation of PWDFT, which targets general-purpose CPUs, used std::map from the C++ Standard Template Library to perform the look-up. This method suffers from poor performance due to its indirect access, especially when the system size increases. Additionally, the CPEs on SW26010Pro are unable to allocate from the main memory shared across cores, making std::map inaccessible on CPEs and hindering pseudopotential lookup acceleration. Considering that there are typically no more than 200 different types of atoms, to address these issues, we replace std::map with a small array. By indexing the array with the atom number, we are able to acquire the pseudopotential with one direct access, resulting in significant performance improvements, as shown in Figure 10. This new mechanism is also more CPE-friendly, as the LD Cache makes loading the pseudopotential efficient in the case of direct indexing. Compared to the previous design, this method boosts the performance of the PW Hamiltonian step by a factor of 10.
Figure 10: Data Structure Optimization of Pseudopotential.
@@ IV-A Setup of the Test Physical Systems and Testing Environment
Figure 11: Testing systems: (a) graphene system and (b) bulk silicon system.
The testing systems include two parts: graphene (labeled as G) and cubic silicon systems (labeled as Si), as shown in Figure 11. These systems are denoted respectively as G11520, G16384, Si64, Si216, Si512, Si1000, Si4096, and Si8192. The corner marks indicate the number of atoms in each system. Notably, our parallelization approach is well-suited for handling large-scale quantum simulations, making it applicable to other complex physical systems such as amorphous silicon [44, 53].
In our experiment, the kinetic energy cutoff $E_{cut}$ is set to 5 Hartree unless otherwise specified. This value has been found sufficient for achieving convergence in silicon crystal calculations [4]. We use the same $E_{cut}$ in graphene systems for consistency. Our optimization approach has good effects for different values of $E_{cut}$. All numerical tests are performed on the new Sunway supercomputer, and each MPI process is applied to one CG (including 1 MPE and 64 CPEs) with separated 16 GB of memory. We consider the time per outer SCF iteration as the time to solution.
@@ IV-B Strong Scalability
The speedup is compared with the minimum number of CGs (64 CGs for Si4096 and 512 CGs for Si8192) capable of performing the calculation. We use 64 to 2048 CGs when simulating Si4096 to test the strong scalability of PWDFT and 512 to 4096 CGs for Si8192.
Figure 12: Strong scaling of PWDFT.
The results are shown in Figure 12. From the figure, we can see that, when scaling up to 2048 CGs (Si8192), PWDFT maintains a parallel efficiency of 48.5%, which is quite acceptable for PW-based DFT calculations. However, the parallel efficiency is not satisfactory when scaling up to half the number of atoms. Even so, we must note that such mediocre parallel efficiency is already much better than the original version and other PW-based DFT software. In Section III-D, due to the poor scalability of Syevd, the optimal process number is always selected in our granularity-aware parallelism for Syevd in the range of $P_{opt}\in[P_{min},2*A]\land P_{opt}=2^{k},k\in\mathbb{N}^{+}$. Hence, when the number of processes reaches a sufficient threshold to achieve the best performance for Syevd, it remains unchanged. Considering that Syevd often occupies a significant portion of the total computation time, it is reasonable to conclude that PWDFT does not exhibit excellent strong scalability and parallel efficiency. Additionally, massive global communication, as discussed in Section II-C2, also partially contributes to this inefficiency.
Figure 13: Strong scaling of PWDFT without syevd.